OpenStudy (anonymous):

" If 3 over x = 7 over y, then ________?"

OpenStudy (anonymous):

3/x = 7/y then x over y = 3 over 7

OpenStudy (maheshmeghwal9):

\[\color{red}{{3 \over x} = {7 \over y}.}\]\[3y=7x.\]\[\color{blue}{\frac{x}{y}=\frac{7}{3}.}\]

OpenStudy (maheshmeghwal9):

Sorry it must be this: -\[\color{blue}{\frac{x}{y}=\frac{3}{7}.}\]

OpenStudy (maheshmeghwal9):